how do you find the perimeter and area of a square whose side is y-3?
: how do you find the perimeter and area of a square whose side is y-3?

Answer by ankor@dixie-net.com(4684)

P = 4(y-3)
P = 4y - 12
:
A = (y-3)^2
FOIL
A = y^2 - 6y + 9
Question 155739: how do you find the perimeter and area of a square whose side is y-3?
: how do you find the perimeter and area of a square whose side is y-3?

Answer by jim_thompson5910(9929)

Perimeter of a Square: P=4s where "P" is the perimeter and "s" is the side length


P=4s Start with the general formula for the perimeter of a square


P=4(y-3) Plug in the given side length of s=y-3


P=4y-12 Distribute


So the perimeter is P=4y-12





Area of a Square: A=s^2 where "A" is the area and "s" is the side length


A=s^2 Start with the general formula for the area of a square


A=(y-3)^2 Plug in the given side length of s=y-3


A=y^2-3y-3y+9 FOIL


A=y^2-6y+9 Combine like terms.


So the area is A=y^2-6y+9